OpenStudy (anonymous):

Diff eq help please! Solve y' = 1 + exp(y) y' = exp(y)

OpenStudy (anonymous):

2 different equations

OpenStudy (tkhunny):

They are separable. What's stopping you from find the antiderivatives? The first: \(\int\dfrac{dy}{1+e^{y}} = dx\) -- Go!

OpenStudy (anonymous):

Ohhhh...my bad lol.

OpenStudy (tkhunny):

Whoops. Right hand side should be \(\int dx\)

OpenStudy (primeralph):